# Tighter Lower Bounds on Aperiodic Ambiguity Function and Their Asymptotic Achievability

### v2

@@ V-A Comparisons with the Existing AF Lower Bound
Example 1: In Fig.1(a), for LAZ with different $Z_{x}$ and $Z_{y}$, with $N=128$ and $M=6$, one can observe that our proposed bounds (under the weight vector $\mathbf{w}^{B,q}$ with $q=\min\left\{{Z_{x},\left\lfloor\frac{\pi}{\gamma}\right\rfloor+1}\right\}$) are tighter than the benchmark bound (59) [5]. As analyzed in Subsection III-A, the configuration of this example satisfies $q<N-\left\lfloor N\sqrt{\text{AFB}_{\text{ref}}}\right\rfloor$, so the consideration of $D_{\text{opt}}$ is unnecessary. Furthermore, the proposed bounds exhibit a corresponding increase as $Z_{x}$ or $Z_{y}$ increases, which is a distinct behavior compared to the benchmark (59). This also indicates that our proposed bounds align more with the physical significance. The rationale behind this is that achieving lower (non-trivial) AF magnitude is increasingly challenging for a larger LAZ. Therefore, it is logical that the AF lower bound should increase with the enlargement of the LAZ.
@@ V-B Asymptotic Achievability of the Proposed Bounds
In Fig. 2, we show the maximum AAF/CAF magnitudes of a Chu sequence pair $\mathbf{s}^{1}$ and $\mathbf{s}^{2}$ with $a_{1}=20$ and $a_{2}=19$ for different values of $N$, along with the proposed aperiodic AF lower bounds associated to the LAZ ($Z_{x}=\left\lfloor\frac{9N}{10|a_{1}|}\right\rfloor$ and $Z_{y}=a_{2}$ according to Corollary 8). We set $\theta_{a}=0.4802\sqrt{\frac{N}{a_{1}}}$ as the asymptotic value of the maximum AAF magnitude for $\mathbf{s}^{1}$, which is smaller than that of $\mathbf{s}^{2}$. According to the tradeoff between $\theta_{a}$ and $\theta_{c}$ as shown in (33a), $\theta_{c}$ is represented by the purple line in the figure.
@@ VI Conclusions
In this work, we have developed new aperiodic AF lower bounds for unimodular sequences with respect to the LAZ in the delay-Doppler plane. We have also presented the trade-offs between the lower bounds for the aperiodic AAF and CAF. Unlike the Levenshtein correlation bound, our first innovation is the introduction of two weight vectors in order to characterize the individual influences of the delay and Doppler shifts, respectively. Moreover, our bounds are derived by leveraging certain insights of the zero delay and $D$ last delays that are associated with the structural properties of aperiodic AF. Existing aperiodic correlation bounds, such as the Welch bound, Sarwate Bound, Levenshtein bounds, Peng-Fan bounds, and Arlery-Tan-Rabaste bounds, can be considered as special cases of our proposed AF lower bounds. Furthermore, it has been shown that our proposed AF lower bounds are asymptotically achievable by the Chu sequence sets and are tighter than the existing benchmark bound.

### v3

@@ V-A Comparisons with the Existing AF Lower Bound
Example 1: In Fig.1(a), for LAZ with different $Z_{x}$ and $Z_{y}$, with $N=128$ and $M=6$, one can observe that our proposed bounds (under the weight vector $\mathbf{w}^{B,q}$ with $q=\min\left\{{Z_{x},\left\lfloor\frac{\pi}{\gamma}\right\rfloor+1}\right\}$) are tighter than the benchmark bound (44) [5]. As analyzed in Subsection III-A, the configuration of this example satisfies $q<N-\left\lfloor N\sqrt{\text{AFB}_{\text{ref}}}\right\rfloor$, hence the consideration of $D_{\text{opt}}$ is unnecessary. Furthermore, the proposed bounds exhibit a corresponding increase as $Z_{x}$ or $Z_{y}$ increases, which is a distinct behavior compared to the benchmark (44). This also indicates that our proposed bounds align more with the physical significance. The rationale behind this is that achieving lower (non-trivial) AF magnitude is increasingly challenging for a larger LAZ. Therefore, it is logical that the AF lower bound should increase with the enlargement of the LAZ.
@@ V-B Asymptotic Achievability of the Proposed Bounds
In Fig. 2, we show the maximum AAF/CAF magnitudes of a Chu sequence pair $\mathbf{s}^{1}$ and $\mathbf{s}^{2}$ with $a_{1}=20$ and $a_{2}=19$ for different values of $N$, along with the proposed aperiodic AF lower bounds associated to the LAZ ($Z_{x}=\left\lfloor\frac{9N}{10|a_{1}|}\right\rfloor$ and $Z_{y}=a_{2}$ according to Corollary 5). We set $\theta_{a}=0.4802\sqrt{\frac{N}{a_{1}}}$ as the asymptotic value of the maximum AAF magnitude for $\mathbf{s}^{1}$, which is smaller than that of $\mathbf{s}^{2}$. According to the tradeoff between $\theta_{a}$ and $\theta_{c}$ as shown in (29a), $\theta_{c}$ is represented by the purple line in the figure. Note that some values in Fig. 2 exceed 00 dB because $\sqrt{N}$ is used as the normalization factor here for calculating decibels instead of $N$.
@@ VI Conclusions
In this work, we have derived generalized Arlery-Tan-Rabaste-Levenshtein AF lower bounds with respect to the LAZ in the delay-Doppler plane. We have also presented the trade-offs between the lower bounds for the aperiodic AAF and CAF. Our first innovation is the introduction of two weight vectors in order to characterize the individual influences of the delay and Doppler shifts, respectively. Moreover, our bounds are derived by exploring certain insights of the zero delay and $D$ last delays that are associated with the structural properties of aperiodic AF. Furthermore, it has been shown that our proposed AF lower bounds are asymptotically achievable by the Chu sequence sets and are tighter than the existing benchmark bound.
Appendix A Proof of Corollary 2
By substituting $\mathbf{w}^{A,q}$ and $D=0$ into (22a), we have
${\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{q}\right)$
$\geq$ $M\left(N-\sum_{s,t=0}^{Z_{x}-1}l_{s,t,N}w_{s}w_{t}-\frac{N^{2}}{MZ_{y}}\sum_{i=0}^{Z_{x}-1}w_{i}^{2}\right)$
$=$ $M\left(N-\sum_{s,t=0}^{q-1}|s-t|\frac{1}{q^{2}}-\frac{N^{2}}{qMZ_{y}}\right)$
$=$ $M\left(N-\frac{(q-1)(q+1)}{3q}-\frac{N^{2}}{qMZ_{y}}\right).$ (47)
Based on ${\theta}_{\text{max}}^{2}\left(M-\frac{1}{q}\right)\geq{\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{q}\right)$, we have ${\theta}_{\text{max}}^{2}\geq\frac{3qMNZ_{y}-3N^{2}-q^{2}MZ_{y}+MZ_{y}}{3(qM-1)Z_{y}}$.
${\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{\sqrt{\frac{3N^{2}}{MZ_{y}}}}\right)$
$\geq$ ${\theta}_{c}^{2}(M-1)+{\theta}_{a}^{2}\left(1-\frac{1}{\hat{q}}\right)$
$\geq$ $M\left(N-\frac{\hat{q}^{2}-1}{3q}-\frac{N^{2}}{\hat{q}MZ_{y}}\right)$
$\geq$ $M\left(N-\frac{2N}{\sqrt{3MZ_{y}}}\right).$ (48)
Appendix B Proof of Theorem 3
Since $A_{\mathbf{s}}(\tau,\nu)={A^{\ast}_{\mathbf{s}}(-\tau,-\nu)}$, we only need to consider $\tau>0$ to analyze the maximum AF magnitude in a given LAZ. Note that if $\nu-a\tau\neq uN$ for any integer $u$, $\sum_{n=0}^{N-1}e^{\frac{j2\pi(\nu-a\tau)n}{N}}=0$ [28]. For $\tau>0$, we have
$\left|A_{\mathbf{s}}(\tau,\nu)\right|$ $=\left|\sum_{n=0}^{N-1-\tau}{s}_{n}^{a}{\left({s}_{n+\tau}^{a}\right)}^{\ast}e^{\frac{j2\pi\nu n}{N}}\right|$
$=\left|\sum_{n=0}^{N-1-\tau}e^{\frac{j\pi an^{2}}{N}}e^{\frac{-j\pi a(n+\tau)^{2}}{N}}e^{\frac{j2\pi\nu n}{N}}\right|$
$=\left|\sum_{n=0}^{N-1-\tau}e^{\frac{j2\pi(\nu-a\tau)n}{N}}\right|$
$=\left|\sum_{l=0}^{N-1-(N-\tau)}e^{\frac{-j2\pi(\nu+a(N-\tau))l}{N}}\right|$
$=\left|A_{\mathbf{s}}(N-\tau,-\nu)\right|.$ (49)
[…]
Case 1: $1\leq\tau\leq 0.45\sqrt{\frac{N}{|a|}}.$
From the choice of $a$, $\nu$ and $\tau$, it is clear that $\nu-a\tau\neq uN$, holds for any integer $u$. By (B), Lemma 2 and for $1\leq\tau\leq 0.45\sqrt{\frac{N}{|a|}}$, we have
$\left|A_{\mathbf{s}}(\tau,\nu)\right|$ $=\left|A_{\mathbf{s}}(N-\tau,-\nu)\right|\leq N-(N-\tau)=\tau$
$\leq 0.45\sqrt{\frac{N}{|a|}}.$
Thus
$\frac{\left|A_{\mathbf{s}}(\tau,\nu)\right|}{\sqrt{N}}\leq\frac{0.45}{\sqrt{|a|}},\quad\text{for Case 1}.$ (50)
Case 2: $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1.
Since $N\geq 5|a|$, we have
$0<-|a|+1+0.45\sqrt{5}|a|$
$\leq-|a|+1+0.45\sqrt{|a|N}<\nu-a\tau<\beta N.$
Note that
$\left|A_{\mathbf{s}}(\tau,\nu)\right|$
$=$ $\left|\sum_{n=0}^{N-1-\tau}e^{\frac{j2\pi(\nu-a\tau)n}{N}}\right|=\left|\frac{1-e^{\frac{j2\pi(\nu-a\tau)(N-\tau)}{N}}}{1-e^{\frac{j2\pi(\nu-a\tau)}{N}}}\right|$
$=$ $\sqrt{\frac{1-\cos\frac{2\pi(\nu-a\tau)(N-\tau)}{N}}{1-\cos\frac{2\pi(\nu-a\tau)}{N}}}=\sqrt{\frac{1-\cos\frac{2\pi(\nu-a\tau)\tau}{N}}{1-\cos\frac{2\pi(\nu-a\tau)}{N}}}$
$=$ $\left|\frac{\sin\frac{\pi(\nu-a\tau)\tau}{N}}{\sin\frac{\pi(\nu-a\tau)}{N}}\right|.$ (51)
Then, we divide Case 2 into the following two sub-cases:
Sub-case 2-1: $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1 with $\nu-a\tau\geq\frac{\beta}{\sin(\beta\pi)}\sqrt{5|a|N}$.
Since $\beta<1$, $\sin x$ is concave downward in $[0,\beta\pi]$. Thus $\sin x>\frac{\sin(\beta\pi)}{\beta\pi}x>0$, for $0<x<\beta\pi$. Since $0<\nu-a\tau<\beta N$, we have $0<\frac{\pi(\nu-a\tau)}{N}<\beta\pi$. Hence
$\sin\frac{\pi(\nu\!-\!a\tau)}{N}\!>\!\frac{\sin(\beta\pi)}{\beta\pi}\!\cdot\!\frac{\pi(\nu\!-\!a\tau)}{N}\!=\!\frac{\sin(\beta\pi)}{\beta N}\!\cdot\!(\nu\!-\!a\tau)\!>\!0.$
Then based on (B), we have
$\frac{\left|A_{\mathbf{s}}(\tau,\nu)\right|}{\sqrt{N}}$
$=$ $\frac{\left|\sin\frac{\pi(\nu-a\tau)\tau}{N}\right|}{\sqrt{N}\left|\sin\frac{\pi(\nu-a\tau)}{N}\right|}<\frac{1}{\sqrt{N}\frac{\sin(\beta\pi)}{\beta N}\cdot(\nu-a\tau)}$
$=$ $\frac{\beta}{\sin(\beta\pi)}\cdot\frac{\sqrt{N}}{(\nu-a\tau)}$
[…]
Sub-case 2-2: $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1 with $\nu-a\tau<\frac{\beta}{\sin(\beta\pi)}\sqrt{5|a|N}$.
In this sub-case, $\frac{\nu-a\tau}{N}\rightarrow 0$ as $N\rightarrow+\infty$. Following (B), for $N\rightarrow+\infty$, we have
$=\frac{N\tau}{\pi(\nu-a\tau)}F\left(\frac{2\pi(\nu-a\tau)\tau}{N}\right),$ (53)
where $F(\varphi)=\frac{1-\cos\varphi}{\varphi}$ with
$\varphi=$ $\frac{2\pi(\nu-a\tau)\tau}{N}=\frac{\tau}{\nu-a\tau}\frac{2\pi(\nu-a\tau)^{2}}{N}$
$<$ $\frac{2\pi\tau}{\nu-a\tau}\cdot\frac{5\beta^{2}|a|}{\sin^{2}(\beta\pi)}.$
For $\varphi>2\pi$, $F(\varphi)\leq\frac{2}{\varphi}\leq\frac{1}{\pi}\approx 0.3183$. For $0<\varphi\leq 2\pi$, the maximum value of $F(\varphi)$ equals $F(\varphi_{0})\approx 0.7246$, $\varphi_{0}\approx 2.3311$ [23].
Then, we consider the $\frac{\tau}{\nu-a\tau}$ part in (B). Note that $\frac{\tau}{\nu-a\tau}=\frac{1}{\frac{\nu}{\tau}-a}$, $10.45\sqrt{\frac{N}{|a|}}<\tau\leq\frac{\beta N}{|a|}-$ 1, and $a<\nu<-a$, we have
$\frac{1}{\frac{-a-1}{0.45\sqrt{N/|a|}}-a}\leq\frac{\tau}{\nu-a\tau}\leq\frac{1}{\frac{a+1}{\beta N/|a|-1}-a}.$
For $N\rightarrow+\infty$,
$\frac{1}{\frac{-a-1}{0.45\sqrt{N/|a|}}-a}\rightarrow\frac{1}{|a|}\quad\text{and}\quad\frac{1}{\frac{a+1}{\beta N/|a|-1}-a}\rightarrow\frac{1}{|a|},$
so $\frac{\tau}{\nu-a\tau}\rightarrow\frac{1}{|a|}$. Thus, from (B), we have
$\lim_{N\to+\infty}\frac{\max_{(\tau,\nu)}{\left|A_{\mathbf{s}}(\tau,\nu)\right|}}{\sqrt{N}}=\sqrt{\frac{F(\varphi_{0})}{\pi|a|}}=\frac{0.4802}{\sqrt{|a|}},$
[…]
Overall, based on the results from (50), (B), and (B), we conclude (40).